OpenStudy (anonymous):

Which of the following is the best thing to do as a first step in using substitution to solve the system of equations? A. Solve the top equation for x. B. Solve the top equation for y. C. Solve the bottom equation for x. D. Solve the bottom equation for y.

OpenStudy (richyw):

it's a bit of topic to show the whole solutions. but I am just trying to explain how it doesn't really matter. This system of equations\[\alpha x+\beta y=\gamma \]\[\zeta x+\xi y=\eta\] Is the same as\[\zeta x+\xi y=\eta\]\[\alpha x+\beta y=\gamma \] where \(\alpha,\beta,\gamma,\zeta,\xi,\eta\) could be any numbers. So how could you possibly know which if the top or the bottom one is correct?

OpenStudy (richyw):

*correct to solve for first

OpenStudy (richyw):

basically just solve which ever one is best for the situation.